OpenStudy (anonymous):

Calculating savings rate If 5000 patients=3.9% savings and 5999=3.6%savings, what is the savings rate for 5456 patients, etc.? How to calculate the numbers in between?

OpenStudy (amistre64):

this is just a linear function since it gives 2 points of reference

OpenStudy (amistre64):

find the slope between the given "points" and determine the linear equation

OpenStudy (amistre64):

x : y --------- 5999 : 3.6 5000 : 3.9 <- subtract --------- 999 : -.3 slope = y/x = -.3/999

OpenStudy (amistre64):

your equation becomes: 3.9 = 5000(-.3/999) + b 3.9 + (1500/999) = b

OpenStudy (amistre64):

y = (-.1/333)x + 500/333 + 39/10 when x = 5456; y = 3.76%

OpenStudy (anonymous):

Thanks a ton! Any advice on how to repeat this multiple times in excel?